# MIGE: A Unified Framework for Multimodal Instruction-Based Image Generation and Editing

### v1

@@ Abstract
Despite significant progress in diffusion-based image generation, subject-driven generation and instruction-based editing remain challenging. Existing methods typically treat them separately, struggling with limited high-quality data and poor generalization. However, both tasks require capturing complex visual variations while maintaining consistency between inputs and outputs. Therefore, we propose MIGE, a unified framework that standardizes task representations using multimodal instructions. It treats subject-driven generation as creation on a blank canvas and instruction-based editing as modification of an existing image, establishing a shared input-output formulation. MIGE introduces a novel multimodal encoder that maps free-form multimodal instructions into a unified vision-language space, integrating visual and semantic features through a feature fusion mechanism. This unification enables joint training of both tasks, providing two key advantages: (1) Cross-Task Enhancement: By leveraging shared visual and semantic representations, joint training improves instruction adherence and visual consistency in both subject-driven generation and instruction-based editing. (2) Generalization: Learning in a unified format facilitates cross-task knowledge transfer, enabling MIGE to generalize to novel compositional tasks, including instruction-based subject-driven editing. Experiments show that MIGE excels in both subject-driven generation and instruction-based editing while setting a state-of-the-art in the new task of instruction-based subject-driven editing. Code and model have been publicly available at this link.
MIGE: A Unified Framework for Multimodal Instruction-Based
Image Generation and Editing
Xueyun Tian♠♡, Wei Li, Bingbing Xu, Yige Yuan♠♡, Yuanzhuo Wang, Huawei Shen♠♡ CAS Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China {tianxueyun23z, xubingbing, yuanyige20z, wangyuanzhuo, shenhuawei}@ict.ac.cn    weili.ucas.ict@gmail.com
Figure 1: Demonstrating the comprehensive capabilities of MIGE. As a unified framework, MIGE excels in subject-driven generation and instruction-based editing while performing well in the new compositional task of instruction-based subject-driven editing.
1 Introduction
@@
Recent advances in diffusion models Rombach et al. (2022); Peebles and Xie (2023); Labs (2023); Esser et al. (2024) have greatly advanced customized image generation. Central to this field are subject-driven image generation, which focuses on preserving a given subject, and instruction-based image editing, which emphasizes flexible modifications based on textual instructions.
@@
Existing methods address these tasks separately, each tailored to a specific input-output format, limiting generalization and hindering a unified instruction-following model. Subject-driven generation aims to preserve subject identity while adhering to textual instructions. Fine-tuning methods, like DreamBooth Ruiz et al. (2023), require separate training for each subject during testing, which limits generalization. In contrast, zero-shot methods like KOSMOS-G Pan et al. (2023) rely on multimodal alignment but struggle with instruction following, mainly due to the scarcity of high-quality multimodal training data, especially for multi-subject inputs. Instruction-based editing, exemplified by InstructPix2Pix Brooks et al. (2023) and EmuEdit Sheynin et al. (2024), modifies existing images based on instructions. These methods require large, high-quality editing datasets to handle diverse instructions but struggle with maintaining local-to-global consistency.
@@
Both tasks face significant challenges due to limited data, which hinders their ability to adapt to diverse subjects or editing requirements. Furthermore, they are typically trained independently, lacking cross-task consistency, which prevents a unified understanding of the tasks. Despite these challenges, the two tasks share a fundamental principle: preserving visual consistency between inputs and outputs while capturing visual variations based on instructions. Additionally, they also follow a similar input-output paradigm, with complementary data focuses. Subject-driven image generation trains models to produce subject-accurate images, while instruction-based editing focuses on manipulating images without altering irrelevant identities. The above limitations as well as the complementary nature underscore the need for a unified framework that combines the strengths of both tasks to achieve improved performance.
@@
Inspired by the above analysis, we propose a unified framework, MIGE, which leverages Multimodal Instructions to unify subject-driven Generation and instruction-based Editing. This vision-language representation enables substituting both entities and entire images in text prompts with their visual counterparts, allowing flexible task and instruction combinations. Structurally, we model subject-driven generation as creating an image on a blank canvas and instruction-based editing as modifying an existing image. This coherent input-output mapping simplifies the process while enabling both tasks to reinforce each other through joint training. Moreover, this integrated approach fosters emergent compositional capabilities beyond what either task can achieve alone.
@@
Building on our unified framework, we address two critical challenges. First, for multimodal instruction encoding, existing methods Li et al. (2023a); Pan et al. (2023); Li et al. (2024b) primarily extract semantic features from images by CLIP vision encoder Radford et al. (2021), which is insufficient for preserving fine-grained subject details. To overcome this limitation, we introduce a multimodal encoder equipped with a feature fusion mechanism. This innovation integrates VAE Kingma (2013) visual features into semantic tokens, effectively capturing both detailed visual information and semantics. Secondly, we enhance compositionality through joint training, facilitating instruction-based subject-driven editing. To further optimize performance in this complex task, we develop a novel data construction pipeline based on a Multimodal Large Language Model (MLLM). This pipeline autonomously generates diverse multimodal instructions and corresponding output images. Moreover, to address the absence of an evaluation benchmark for this new composition task, we introduce MIGEBench. This specialized benchmark evaluates compositionality in terms of subject preservation and instruction adherence, providing a comprehensive assessment of the new task.
@@
Our experiments demonstrate that joint training under the MIGE framework enables mutual reinforcement between subject-driven generation and instruction-based editing by leveraging complementary data, leading to significant performance gains over training on individual tasks. Specifically, we observe improved instruction adherence on DreamBench Ruiz et al. (2023) and enhanced detail preservation on EmuEdit and MagicBrush Zhang et al. (2024) test sets. Furthermore, MIGE achieves state-of-the-art results on MIGEBench, showcasing its fine-grained controllability in the emerging compositional tasks. Extensive case studies (as shown in Figure 5 and Figure 6) further illustrate the ability of MIGE to generate precise and consistent outputs, highlighting its effectiveness in diverse scenarios.
@@
  • Unified Framework: We propose MIGE, a unified framework that integrates subject-driven generation and instruction-based editing, facilitating joint training and mutual enhancement of both tasks.
  • Compositional Capability: We observed that joint training unlocks the compositional capability of instruction-based subject-driven editing. We propose a novel data construction pipeline and introduce the MIGEBench benchmark for comprehensive evaluation.
  • Strong Performance: The extensive experimental results show that MIGE achieves competitive results in both subject-driven generation and instruction-based editing while establishing a new state-of-the-art in instruction-based subject-driven image editing.
2 Related Work
In this section, we focus on discussing the most relevant studies in this section. A more comprehensive discussion, including additional related work, can be found in the Appendix D.
@@ 2 Related Work
Unifying tasks in image generation is challenging, and several approaches have been proposed. Pixwizard Lin et al. (2024) leverages task vectors to integrate image generation and editing but struggles with vision-language instructions due to architectural limitations. Instruct-Imagen Hu et al. (2024) unifies tasks through multimodal instructions but remains ineffective for editing even after fine-tuning. ACE Han et al. (2024) encodes various inputs as condition units within a transformer but is limited to editing. OmniGen Xiao et al. (2024b) concatenates conditions before the noise and uses autoregressive transformers, but fails to distinguish between background and reference images, resulting in poor background preservation. UniReal Chen et al. (2024) makes finer distinctions but requires complex prompts for task differentiation. In contrast, MIGE ensures input-output consistency with conditional input and multimodal instructions, simplifying task distinction and offering better scalability without the need for complex hierarchical prompts.
@@
Subject-driven image editing allows image modifications based on a user-specified subject, typically through addition or replacement. Methods like DreamEdit Li et al. (2023b) and DreamMix Yang et al. (2024) require multiple input images for fine-tuning, associating new concepts with special tokens, followed by iterative inpainting. PBE Yang et al. (2023) conditions a diffusion model on compressed reference image representations, while TIGIC Li et al. (2025) integrates reference images by positioning them in predefined backgrounds. MADD He et al. (2024) offers more diverse control by providing masks, boxes, or target coordinates. Overall, these methods rely on multiple reference images or precise cues like bounding boxes and masks, and cannot follow raw textual instructions for accurate and controllable generation, making them inflexible.
Figure 2: Demonstration of MIGE as a unified framework for processing multimodal instructions and conditional inputs across diverse tasks and scenarios.
Figure 3: Overall framework of MIGE. MIGE consists of two components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model for modeling input-output relationships. The encoder incorporates a feature fusion mechanism to integrate visual and semantic features from reference image.
3 Unified Framework
3.1 Task Unification
@@
Unified Multimodal Instruction
@@
To enable joint training across multiple tasks, a unified task representation is essential. We introduce multimodal instructions composed of interleaved images and text, which provide both visual references and textual guidance for various controllable generation tasks. As shown in Figure 2, “<imagehere>” serves as a placeholder, sequentially replaced by input images, which can be the reference subject or the entire scene, complementing the semantics to form interleaved expressions. This unified approach effectively accommodates subject-driven generation and instruction-based editing, while remaining extensible to the more complex combination task.
Unified Conditional Input
We adopt a conditional input design to unify tasks structurally, ensuring clear task distinction while enabling shared capabilities. By concatenating different conditions, we differentiate initial generation states and capture task-specific nuances, enhancing execution accuracy. As shown in Figure 2, for instruction-based editing, we concatenate the VAE-encoded source image with the noise tensor, guiding generation based on the given image. For subject-driven generation, we use an all-zero tensor akin to a blank canvas, prompting the model to generate freely while preserving specified visual features. This design effectively captures the differences between the two tasks while ensuring input-output consistency in editing. Moreover, structuring tasks within a unified framework allows the model to leverage shared capabilities and easily extend to new tasks.
3.2 Architecture
((a))
((b))
Figure 4: Data construction pipelines for instruction-based subject-driven image editing.
@@ 3.2 Architecture
The architecture of MIGE, as shown in Figure 3, consists of two main components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model Peebles and Xie (2023) for modeling input-output relationships. The diffusion model takes concatenated latent noise and conditional input along the channel dimension as input, performing controllable generation under the control of multimodal conditions. To further enhance the integration of visual and semantic information from the reference image, we introduce a novel feature fusion mechanism in the encoder.
Multimodal Encoder
@@
To map the multimodal instructions into a unified vision-language semantic space, we design a multimodal encoder comprising a Large Language Model (LLM) and an image feature encoding component, which includes a pretrained VAE encoder Kingma (2013) for visual feature extraction, a pretrained ViT from EVA-CLIP Fang et al. (2023) for semantic feature extraction, a Q-Former Li et al. (2023a), and a linear projection layer. Each image is represented by 32 tokens, which are then encoded by the LLM alongside text tokens to form a unified multimodal condition.
@@
Unlike prior works Li et al. (2024a); Pan et al. (2023); Li et al. (2024b) that mainly extract semantic features of the reference images, which lack the fine-grained details necessary for preserving subject-specific features. To address this, we propose a feature fusion mechanism that combines the strengths of different visual encoders. ViT serves as a semantic feature extractor, while a VAE encoder acts as a visual feature extractor, leveraging its ability to compress and reconstruct images. As shown in Figure 3, we use the ViT semantic features compressed by the Q-Former, denoted as $f_{s}$, as guidance to adaptively incorporate the visual features $f_{v}$ extracted by the VAE. The fusion mechanism is expressed as: $f_{\text{img}}=f_{s}+\text{MLP}(\text{Attn}(f_{s},f_{v}))$ where $\text{Attn}(f_{s},f_{v})=\frac{Q(f_{s})\cdot K(f_{v})}{\sqrt{d}}V(f_{v})$. Here $d$ denotes the output dimension of the features, $Q$ is a linear layer, and both $K$ and $V$ are two-layer networks, where $K(\cdot)=V(\cdot)=\text{Linear}(\text{GELU}(\text{Linear}(\cdot)))$.
Through this fusion mechanism, we can effectively capture both visual and semantic information of reference images simultaneously without introducing additional image tokens.
3.3 Joint Training

### v3

# MIGE: Mutually Enhanced Multimodal Instruction-Based Image Generation and Editing

@@ Abstract.
Despite significant progress in diffusion-based image generation, subject-driven generation and instruction-based editing remain challenging. Existing methods typically treat them separately, struggling with limited high-quality data and poor generalization. However, both tasks require capturing complex visual variations while maintaining consistency between inputs and outputs. Inspired by this, we propose MIGE, a unified framework that standardizes task representations using multimodal instructions. It first treats subject-driven generation as creation on a blank canvas and instruction-based editing as modification of an existing image, establishing a shared input-output formulation, then introduces a novel multimodal encoder that maps free-form multimodal instructions into a unified vision-language space, integrating visual and semantic features through a feature fusion mechanism. This unification enables joint training of both tasks, providing two key advantages: (1) Cross-Task Enhancement: by leveraging shared visual and semantic representations, joint training improves instruction adherence and visual consistency in both subject-driven generation and instruction-based editing. (2) Generalization: learning in a unified format facilitates cross-task knowledge transfer, enabling MIGE to generalize to novel compositional tasks, including instruction-based subject-driven editing. Experiments show that MIGE excels in both subject-driven generation and instruction-based editing while setting a SOTA in the new task of instruction-based subject-driven editing. Code and model have been publicly available at this link.
Controllable Image Generation, Diffusion Model, Joint Training
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33th ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06submissionid: 6167ccs: Computing methodologies Image representations
1. Introduction
@@
Recent advances in diffusion models(Rombach et al., 2022; Peebles and Xie, 2023; Labs, 2023; Esser et al., 2024) have significantly advanced the development of customized image generation, enabling its widespread adoption in both subject-driven generation(Ruiz et al., 2023; Ye et al., 2023; Li et al., 2024a; Pan et al., 2023; Kang et al., 2025; Kim et al., 2024; Wu et al., 2025; Liang et al., 2025; Peng et al., 2024) and instruction-based editing(Brooks et al., 2023; Sheynin et al., 2024; Fu et al., 2023; Han et al., 2024). The former focuses on preserving a given subject, while the latter emphasizes flexible modifications guided by textual instructions. Their capability to generate personalized and controllable results demonstrates substantial practical value, thereby attracting growing research attention.
@@
Existing approaches (Li et al., 2024a; Pan et al., 2023; Hu et al., 2024; Huang et al., 2024b) typically treat these two tasks independently and address them separately by tailoring specific input-output formats to each. However, this results in limited generalization capability and necessitates extensive training data. Specifically, subject-driven generation aims to preserve subject identity while adhering to textual instructions. Fine-tuning methods, like DreamBooth (Ruiz et al., 2023), require separate training for each subject during testing, which limits generalization. In contrast, zero-shot methods like KOSMOS-G (Pan et al., 2023) rely on multimodal alignment but struggle with instruction following, mainly due to the scarcity of high-quality multimodal training data, especially for multi-subject inputs. Instruction-based editing, exemplified by InstructPix2Pix (Brooks et al., 2023), modifies images based on instructions. These methods require large, high-quality editing datasets to handle diverse instructions but struggle with maintaining local-to-global consistency.
@@
The above separate processing methods can only rely on limited data and single-task optimization objectives, which hinders their ability to adapt to diverse subjects or editing requirements, and lack cross-task consistency, preventing a unified understanding of the tasks. To address these challenges, we shift our perspective from separate processing to the objectives of these two tasks. These two tasks share a fundamental principle: maintaining visual consistency between inputs and outputs while capturing variations guided by instructions. They also follow a similar input-output paradigm with complementarities: subject-driven generation emphasizes subject accuracy, while instruction-based editing manipulates images without altering irrelevant identities. These limitations and complementarities highlight the need for a unified framework that takes advantage of the strengths of both tasks to improve performance.
@@
Inspired by the above analysis, we propose a unified framework, MIGE, which leverages Multimodal Instructions to unify subject-driven image Generation and instruction-based Editing. The vision-language representation enables substituting both entities and entire images in text prompts with their visual counterparts, allowing flexible task and instruction combinations. Structurally, we model subject-driven generation as creating an image on a blank canvas and instruction-based editing as modifying an existing image. This coherent input-output mapping simplifies the process while enabling both tasks to reinforce each other through joint training. Moreover, this integrated approach fosters emergent compositional capabilities beyond what either task can achieve alone.
@@
Building on our unified framework, we address two critical challenges. First, for multimodal instruction encoding, existing methods (Li et al., 2023b; Pan et al., 2023; Li et al., 2024c) mainly rely on CLIP vision encoders (Radford et al., 2021) to extract semantic features, which is insufficient for preserving fine-grained subject details. To overcome this limitation, we propose a multimodal encoder with a feature fusion mechanism that integrates VAE (Kingma, 2013) visual features into semantic tokens, effectively capturing both detailed visual information and semantics. Secondly, we enhance compositionality through joint training, facilitating instruction-based subject-driven editing. To further optimize performance in this complex task, we develop a novel data construction pipeline based on a Multimodal Large Language Model (MLLM). This pipeline autonomously generates diverse multimodal instructions and corresponding output images. Moreover, to address the absence of an evaluation benchmark for this new composition task, we introduce MIGEBench. This specialized benchmark evaluates compositionality in terms of subject preservation and instruction adherence, providing a comprehensive assessment of the new task.
@@
Our experiments demonstrate that joint training under the MIGE framework enables mutual enhancement between subject-driven generation and instruction-based editing, leading to significant performance gains over training on individual tasks. As shown in Figure 2, joint training markedly improves input-output consistency across both tasks, especially in editing, with scores rising from 0.821 to 0.873. Moreover, MIGE achieves state-of-the-art results on MIGEBench, showcasing its fine-grained controllability in emerging compositional tasks. Extensive case studies (Figures 78) further demonstrate the ability of MIGE to generate precise and consistent outputs, highlighting its effectiveness across diverse scenarios.
Figure 2. Joint training significantly improves input-output similarity across both tasks. We show the average DINO and CLIP-I scores for subject-driven generation and instruction-based editing under single-task and joint training.
@@
  • Promising perspective of task mutual enhancement: We shift our perspective from traditional separate processing to mutual enhancement. Then, we design a unified input-output format for joint training on subject-driven generation and instruction-based editing, validating that joint training enables task enhancement.
  • Novel framework for Compositional Capability: We propose MIGE, a unified framework integrating two mainstream tasks in controllable image generation. Joint training further enables the compositional capability of instruction-based subject-driven editing. We also introduce a novel data pipeline and MIGEBench for comprehensive evaluation.
  • Strong Performance with Extensive Experiments: Extensive results show that MIGE performs competitively on subject-driven generation and instruction-based editing, with scores rising from 0.821 to 0.873, and sets a new state-of-the-art in instruction-based subject-driven editing.
2. Related Work
In this section, we discuss the most relevant studies. A more comprehensive discussion can be found in the Supplementary Material.
@@ 2. Related Work
Unifying tasks in image generation remains challenging. Some methods (Li et al., 2025a; Huang et al., 2024a) focus on bridging image understanding and generation, such as SEED-X (Ge et al., 2024), Emu2 (Sun et al., 2024), and WeGen (Huang et al., 2025), which use autoregressive models with auxiliary diffusion modules. UnifiedMLLM (Li et al., 2024b) employs MLLMs to learn shared representations and route to task-specific models. However, these approaches either emphasize understanding over generation or rely on complex pipelines. Others aim to unify generation and editing. Pixwizard (Lin et al., 2024) uses text-only instructions and task vectors. Instruct-Imagen (Hu et al., 2024) supports multimodal inputs but performs poorly on editing. ACE (Han et al., 2024) is limited to editing via transformer-based conditioning. OmniGen (Xiao et al., 2024) cannot distinguish reference images. OmniControl (Tan et al., 2024) merges all conditions with noise, while UniReal (Chen et al., 2024c) adds finer control but requires complex prompts. RealGeneral (Lin et al., 2025) enhances attention, and OnePic (Tao et al., 2024) depends on multiple references and masks. Overall, prior work often sacrifices flexibility or generality, combining tasks without leveraging their intrinsic connections. In contrast, we propose MIGE, which jointly trains on subject-driven generation and instruction-based editing, revealing their mutual benefits.
@@
Subject-driven image editing enables modifications based on a user-specified subject, typically including addition or replacement. Existing methods often rely on multiple reference images or precise cues. DreamEdit (Li et al., 2023a) and DreamMix (Yang et al., 2024) require multiple images for fine-tuning and iterative inpainting. PBE (Yang et al., 2023) uses compressed reference embeddings, while TIGIC (Li et al., 2025c) embeds subjects into predefined backgrounds. MADD (He et al., 2024a) supports diverse controls via masks or coordinates. AnyDoor (Chen et al., 2024b), AnyEdit (Yu et al., 2024), MimicBrush (Chen et al., 2024a) and BlobCtrl (Li et al., 2025b) require masks or additional range blobs. FreeEdit (He et al., 2024b) is limited to instruction templates and class-aware features. Overall, these approaches depend on detailed inputs and lack flexibility in handling diverse instructions for controllable generation.
Figure 3. Demonstration of MIGE as a unified framework for processing multimodal instructions and conditional inputs across diverse scenarios.
Figure 4. Overall framework of MIGE. MIGE consists of two components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model for modeling input-output relationships. The encoder incorporates a feature fusion mechanism to integrate visual and semantic features from reference image.
3. Method
In this section, we first shift our view from separate training to mutual enhancement. Guided by this new perspective, we first introduce the unified framework MIGE, followed by details of the architecture design and joint training approach.
3.1. Unified framework MIGE
@@
3.1.1. Unified Multimodal Instruction
@@
To enable joint training across multiple tasks, a unified task representation is essential. We introduce multimodal instructions composed of interleaved images and text, which provide both visual references and textual guidance for various controllable generation tasks. As shown in Figure 3, “¡imagehere¿” serves as a placeholder, sequentially replaced by input images, which can be the reference subject or the entire scene, complementing the semantics to form interleaved expressions. This unified approach effectively accommodates subject-driven generation and instruction-based editing while remaining extensible to the more complex combination task.
3.1.2. Unified Conditional Input
We design a unified conditional input format to distinguish tasks while enabling shared capabilities. As shown in Figure 3, conditional inputs vary across tasks. For instruction-based editing, the VAE-encoded source image is used to preserve fine details of the original image, inform the model to edit only the target region, as done in prior editing methods (Brooks et al., 2023; Zhang et al., 2024). To standardize the input format, we concatenate an all-zero tensor for subject-driven generation to indicate generation from scratch with visual constraints from the multimodal instruction. This design unifies the input format, ensures input-output consistency, and facilitates extensibility to new tasks within a unified framework.
3.2. Architecture
@@
The architecture of MIGE, as shown in Figure 4, consists of two main components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model (Peebles and Xie, 2023) for modeling input-output relationships. The diffusion model takes concatenated latent noise and conditional input along the channel dimension as input, performing controllable generation under the control of multimodal conditions. To further enhance the integration of visual and semantic information from the reference image, we introduce a novel feature fusion mechanism in the encoder.
3.2.1. Multimodal Encoder
@@
To map the multimodal instructions into a unified vision-language semantic space, we design a multimodal encoder comprising a Large Language Model (LLM) and an image feature encoding component, which includes a pretrained VAE encoder (Kingma, 2013) for visual feature extraction, a pretrained ViT from EVA-CLIP (Fang et al., 2023) for semantic feature extraction, a Q-Former (Li et al., 2023b), and a linear projection layer. Each image is represented by 32 tokens, which are then encoded by the LLM alongside text tokens to form a unified multimodal condition.
@@
Prior works (Li et al., 2024a; Pan et al., 2023; Li et al., 2024c) focus on extracting semantics from reference images but lack the fine-grained details needed to preserve subject-specific features. To address this, we propose a feature fusion mechanism that combines the strengths of different visual encoders. ViT serves as a semantic feature extractor, while a VAE encoder acts as a visual feature extractor, leveraging its ability to compress and reconstruct images. As shown in Figure 4, we use the ViT semantic features compressed by the Q-Former, denoted as $f_{s}$, as guidance to adaptively incorporate the visual features $f_{v}$ extracted by the VAE. Inspired by TokenPacker (Li et al., 2024d), we represent rich features with fewer tokens, and apply ReLU (Glorot et al., 2011) to enhance stability. The fusion mechanism is expressed as: $f_{\text{img}}=f_{s}+\text{MLP}(\text{Attn}(f_{s},f_{v}))$ where $\text{Attn}(f_{s},f_{v})=\frac{Q(f_{s})\cdot K(f_{v})}{\sqrt{d}}V(f_{v})$. Here $d$ denotes the output dimension of the features, $Q$ is a linear layer, and both $K$ and $V$ are two-layer networks, where $K(\cdot)=V(\cdot)=\text{Linear}(\text{GELU}(\text{Linear}(\cdot)))$.
This fusion mechanism captures both semantic and visual information without introducing extra image tokens. The multimodal encoder compresses each image into 32 fused tokens, which replace the “¡imagehere¿” placeholder in the interleaved instruction. The LLM then encodes the entire prompt into a unified vision-language representation to guide the diffusion model.
(a) Subject Addition Dataset Construction Pipeline
(b) Subject Replacement Dataset Construction Pipeline
Figure 5. Data construction pipelines for instruction-based subject-driven image editing.
3.3. Joint Training